OpenStudy (anonymous):

Your family is driving on vacation. You have been driving for 3 1/4 hours. You have another 2 3/4 hours before you arrive at your destination. Which equation models the total driving time?

OpenStudy (anonymous):

You have to drive for a total of 6 hours. Talk about a stay-cation.

OpenStudy (anonymous):

i think you have to add them

OpenStudy (anonymous):

What are they going to Wisconsin pffff

OpenStudy (anonymous):

Seriously nowhere good is within 6 hours drive of here

OpenStudy (anonymous):

You have to go far far away

OpenStudy (anonymous):

\[A.) \frac{ 8 }{ 4 } - \frac{ 9 }{ 4 }= B.) \frac{ 13 }{ 4 } - \frac{ 11 }{ 4 }= C.) \frac{ 8 }{ 4 } + \frac{ 9 }{ 4 }= D.) \frac{ 13 }{ 4 } + \frac{ 11 }{ 4 }=?\]

OpenStudy (anonymous):

Satellite is right. Think of it this way, 1/4 hour = 15 minutes 1 hour = 60 minutes

OpenStudy (anonymous):

You have to convert the mixed fraction into an improper fraction. So, 3*4 +1 = 13/4 and 2*4+3 = 11/4. And then you add them together so D is the answer